Show that the function f:[2,3]→R defined by :
f(x)={0 if x is rational {1 if x is irrational
is discontinues and not integrable over [2,3].Does it imply that every discontinues function is non-integrable?Justify your answer.
1
Expert's answer
2019-03-02T16:29:41-0500

Since "\mathbb{Q}" is dense in "\mathbb{R}", we have that "\sup_{x\in [x_i, x_{i+1}]}f(x)=1" and "\inf_{x\in [x_i, x_{i+1}]}f(x)=0" on any interval "[x_i, x_{i+1}] \subset \mathbb{R}", therefore the upper and lower sums don't converge to the same limit. Hence "f" is not Riemann-integrable.


It does not imply that every discontinues function is non-integrable because "g(x)=x" for "x\in [0,1)" and "g(x)=x+1" for "x\in [1,2]" is integrable on "[0,2]" and have discontinuity at "x=1".